OpenStudy (anonymous):

How do you express x³+7x²+10x in factored form

OpenStudy (anonymous):

first take out x: = x(x^2 + 7x + 10) now try to factor whats in the parentheses

OpenStudy (kira_yamato):



OpenStudy (kira_yamato):

x³+7x²+10x = x(x²+7x+10) = x(x+2)(x+5)